# Balancing Embedding Spectrum for Recommendation

### v1

@@ V-C1 Can DirectSpec Prevent Collapse?
Figure 9: (a) shows how the accuracy changes with $\tau_{1}$, (b) displays the sensitivity of $k$, and (c) illustrates how the accuracy changes with $\tau_{0}$ on Yelp.
TABLE VI: Comparison between static and dynamic temperature design.
@@
V-C2 Adaptive Temperature Design
We first compare the dynamic and static designs proposed in Section IV-C, and report the results in Table VI. We observe that the static design outperforms the dynamic design by a large margin on all datasets, and the reason might be attributed to the data scarcity on CF task, making the algorithm difficult to learn the inherent user/item relations from the interactions data in a dynamic way. Therefore, DirectSpec+ is implemented with a static temperature design, which avoids bringing additional complexity. We first tune $\tau_{1}$ by fixing $\tau_{0}=\tau_{1}$ and report how the accuracy changes with $\tau_{1}$ in Figure 9 (a) and (b). The accuracy first increases and then drops as $\tau_{1}$ increases, which is maximized at $\tau_{1}=4$ on Yelp. $\tau_{1}$ controls the strength of penalties on highly correlated unobserved samples (i.e., hard negatives), the samples that are not highly correlated tend to be ignored when $\tau_{1}$ is set too large, while a too small $\tau_{1}$ fails to optimize the highly correlated samples in an effective way. On the other hand, among all unobserved interactions, some are correlated to some extent that should be decorrelated with a slower pace which we use an another hyperparameter $\tau_{0}$ to control. In Figure 9 (b), we observe that optimizing the user/item pairs whose graph distance are $K=2$ with a slower pace $\tau_{0}$ results in a better accuracy, and the accuracy drops after introducing the pairs farther away (i.e., $K>2$) as they are too far away on the graph that should be considered irrelevant. In Figure 9 (c), DirectSpec+ achieves the best performance at $\tau_{0}=3.5$ which is slightly smaller than $\tau_{1}=4$ on Yelp (see Figure 9 (a)), and further decreasing $\tau_{0}=3.5$ leads to a worse performance. This finding indicates that the correlation between the pairs that are close on the graph are still significantly smaller than the observed interactions.
V-C3 Impact of $\alpha$
@@
According to Equation (13), $\alpha$ controls the extent of spectrum balancing. A larger (smaller) $\alpha$ implies that large singular values are multiplied by smaller (larger) weights, leading to a more (less) uniform spectrum distribution. Preposition 1 does not hold when $\alpha$ is too large: $01-\alpha\sigma_{1}^{2}<$ 0 (with $L=1$). Since it takes additional time to calculate $\sigma_{1}$ during each training, we use the grid search for optimal $\alpha$. As shown in Figure 10 (a) and (b), both of the accuracy and erank first increase and then drop, showing a similar trend. Specifically, the speed of collapse tends to outrun the speed of spectrum balancing when $\alpha$ is small, and the drop in erank as increasing $\alpha$ demonstrates our analysis in Section IV-A that a large $\alpha$ instead results in a sharp distribution. In addition, the best performance is achieved at $\alpha=1.1$ on CiteULike and $\alpha=0.6$ on Yelp, and the erank and accuracy are more sensitive when $\alpha$ is small on Yelp, which might be related to the fact that Yelp suffers more from the collapse issue than other datasets.
(a) CiteULike
(b) Yelp
Figure 10: Impact of $\alpha$ on accuracy (nDCG@10) and erank.
Figure 11: Correlations of users (left) and items (right) during the training.
V-C4 User/Item Correlation
As analyzed in Section IV-C, the user/item correlation can reflect the extent of embedding collapse, and the correlation here is measured by Equation (14). As shown in Figure 11, the correlation is low before training, while users/items become highly correlated immediately after training starts on LightGCN(pos) which only optimizes positive (i.e., observed) pairs, causing user and items to have similar representations and resulting in a complete collapse. For LightGCN using BPR loss, the correlation first immediately increases and then gradually drops, whereas the correlation keeps low throughout the training on DirectSpec+, demonstrating its effectiveness to tackle embedding collapse. On the other hand, correlation can also measure the extent of over-smoothing that a higher value indicates more similar representations for GNN-based methods. Since it has been shown that LightGCN helps alleviate the over-smoothing, the lower correlation of DirectSpec+ than LightGCN also indicates the superiority of our proposed method addressing over-smoothing. We also notice that the correlation did not converges to 0, due to the reason that $\mathbf{H}_{U}\mathbf{H}_{U}^{T}$ and $\mathbf{H}_{I}\mathbf{H}_{I}^{T}$ are full rank matrices when they are optimized to identity matrices (i.e., completely decorrelated) according to Equation (14), while the embedding matrix is merely a low-rank approximation: $d\ll min(|\mathcal{U}|,|\mathcal{I}|)$, thus the correlation would not converge to 0 unless $d\geq{\rm\max}(|\mathcal{U}|,|\mathcal{I}|)$.
VI Related Work
VI-A Collaborative Filtering
@@
Collaborative Filtering (CF), a fundamental task for recommender systems, makes predictions based on user’s past behaviours. Early CF methods exploit users that share similar interests or items that tend to be interacted by similar users to infer user preference [34]. Nowadays, model-based methods are prevalent in CF and recommender systems. The core idea is to map the high dimensional sparse recommendation data to a low-dimensional space, where users and items are characterized as learnable vectors and are optimized based on observed interactions [4]. With the development of computer hardware and communication network, more information can be collected to analyze user behaviour with advanced algorithms. Besides user-item interactions, side information such as social relations [3], reviews [35], geological information [36, 37], and so on can help the algorithms better understand user taste and preference. On another line, early simple CF methods [4, 38] are replaced by more powerful algorithms such as neural network [39, 5], attention mechanism [40, 6], transformer [8], etc. Recently, graph neural network (GNN) has also shown tremendous success in recommender systems [15, 25, 11]. By representing interactions as a bipartite graph, the core idea of GNN is to exploit higher-order neighbor connections to facilitate user/item representations [18]. However, it has been reported that multi-layer GNNs suffer from an over-smoothing issue that causes user/item representations to be indistinguishable [31, 19]. We showed that an embedding collapse issue resembling the over-smoothing exists on recommendation algorithms in general, and our proposed approach has also been shown effective tackling over-smoothing.
VI-B Collapse in Representation Learning
@@
Representation learning has attracted tremendous attention in various research fields [41, 42], the goal of which is to represent data in an effective and efficient way to make it easier to extract useful information when building predictors. However, it has also been shown that the model tends to collapse where all inputs are mapped to the same constant vector, when only optimizing the model based on the positive pairs [43]. This issue can be well alleviated by self-supervised learning and contrastive learning by exploiting negative samples in an effective way [44]. Due to the heavy computation cost, some research effort has been made to simplify the algorithms without explicitly sampling negative data [43, 45, 46]. Unfortunately, subsequent works show that a dimensional collapse cannot be ignored where the embedding vectors end up spanning a lower dimensional subspace of the whole embedding space [47, 48]. The collapse issue in representation learning shares similarities to some issues such as over-smoothing in GNN [19], which inspires some researchers to tackle issues in GNN [49, 50].
Inspired by the aforementioned works tackling collapse in representation learning, we reviewed existing recommendation methods, and showed that they suffer from a collapse issue as well. Particularly, the representations tend to collapse to a constant vector when only optimization observed interactions, and an incomplete collapse still exists despite introducing negative samples [12, 5] where the representations are predominantly distributed in certain dimensions. We showed that CL-based methods help tackle this issue to some extent due to the uniformity loss which implicitly balances the embedding spectrum [21, 23, 30]. We address this issue from a spectral perspective by directly balancing the embedding spectrum, which has been demonstrated more effective than existing works. Lastly, we notice a recent work [32] also points out a dimensional collapse issue sharing similarity with our work, while we demonstrated the superiority over it in terms of performance and the effectiveness tackling the collapse issue.
VII Conclusion
@@
In this work, we showed that existing CF methods mostly suffer from an embedding collapse issue. Particularly, optimization solely on the observed interactions causes the representations to collapse to a constant vector, while negative sampling that can decelerate the collapse by balancing the embedding spectrum still results in an incomplete collapse. To tackle this issue, we proposed a DirectSpec which acts as an all pass filter to directly balances the spectrum distribution to ensure that all dimensions can contribute to the user/item embeddings as equally as possible. We conducted comprehensive analysis on DirectSpec from a decorrelation perspective and further proposed an enhanced variant DirectSpec+ to efficiently penalize irrelevant samples with self-paced gradients. In addition, we shed light on the close connection between DirectSpec+ and contrastive learning, that uniformity, a key design contributing to recommendation, can alleviate collapse issue by indirectly balancing spectrum distribution, that can be considered as a special case of DirectSpec+. Finally, extensive experiments on three publicly available datasets demonstrated the effectiveness and efficiency of our proposed methods.

### v2

@@ 5.2.5. Comparison with Solutions to Overs-Smoothing
Figure 10. (a) and (b) show how the accuracy changes with $\tau_{1}$, (c) displays the sensitivity of $K$, and (d) illustrates how the accuracy changes with $\tau_{0}$.
5.3. Model Analysis
5.3.1. Can DirectSpec Prevent Collapse?
In Figure 8 (a), we report how the erank changes as training proceeds on BPR, LightGCN, and DirectSpec+(MF). For BPR and LightGCN, the erank plunges significantly in the first place, then increases slowly and tends to converge to a value smaller than $d$, resulting in an incomplete embedding collapse. While the erank of DirectSpec+(MF) is close to $d$ at the beginning of training and remains almost unchanged throughout the training. Figure 8 (b) illustrates the normalized singular value distribution of the embedding matrix when the accuracy of BPR and DirectSpec+(MF) is maximized. We can see that most singular values of BPR’s embeddings are less than 0.2, barely contributing to the representations; while DirectSpec+(MF)’s singular values are mostly larger than 0.5, indicating that the representations are contributed by more dimensions.
5.3.2. DirectSpec vs. DirectSpec+
The goal of DirectSpec+ is not simply to generate a perfectly balanced spectrum like DirectSpec, but to balance the spectrum while also preserving the mapped relationships among users/items in the embedding space. In Figure 8 (c), the average similarity on pairs in test sets are defined as: $SIM=\frac{\sum_{u}\frac{\sum_{i\in\mathcal{N}^{t}_{u}}\sigma(\mathbf{h}_{u}^{T}\mathbf{h}_{i})}{|\mathcal{N}^{t}_{u}|}}{|\mathcal{U}|}$ where $\mathcal{N}^{t}_{u}$ is the item set $u$ has interacted with in test sets and $\sigma(\cdot)$ is the sigmoid function. In spite of a higher erank DirectSpec has shown during training, DirectSpec+ still outperforms DirectSpec as it shows a higher average similarity on test sets, indicating that DirectSpec+ is better at representing the user/item relationships in the embedding space in a reasonable way.
Table 5. Comparison between static and dynamic temperature design.
@@ 5.3.1. Can DirectSpec Prevent Collapse?
5.3.3. Impact of $\alpha$
@@
$\alpha$ serves as a global coefficient controlling the extent of spectrum balancing. A larger (smaller) $\alpha$ implies that large singular values are multiplied by smaller (larger) weights, leading to a more (less) uniform spectrum distribution. According to Equation (18), DirectSpec could fail when $\alpha$ is set too large: $01-\alpha\sigma_{i}^{2}<$ 0 (with $K=1$). Since it takes additional time to calculate $\sigma_{1}$ during each training, we use the grid search for optimal $\alpha$. As shown in Figure 9, both the accuracy and erank first increase and then drop, showing a similar trend. Specifically, the speed of collapse tends to outrun the speed of spectrum balancing when $\alpha$ is small as erank converges to a value much smaller than $d$. While when $\alpha$ is large, the large singular values in the original distribution are multiplied by too small weights, which might cause a sharp singular value distribution similar to the original one, explaining why erank declines as consistently increasing $\alpha$. In addition, the best performance is achieved at $\alpha=1.1$ on CiteULike and $\alpha=0.6$ on Yelp, and the erank and accuracy are more sensitive when $\alpha$ is relatively small on Yelp, which might be related to the fact that Yelp suffers more from the collapse issue than other datasets.
5.3.4. Adaptive Temperature Design
We first compare the dynamic and static designs proposed in Section 4.3, and report the results in Table 5. We observe that the static design outperforms the dynamic design by a large margin on all datasets, and the reason might be attributed to the data scarcity on CF task, making the algorithm difficult to learn the inherent user/item relations from the interactions data in a dynamic way. Therefore, DirectSpec+ is implemented with a static temperature design, which avoids bringing additional complexity.
For the static temperature design, we first fixed $\tau_{0}=\tau_{1}$ and only tune $\tau_{1}$, then we gradually reduce $\tau_{0}<\tau_{1}$ and observe how accuracy changes accordingly. As shown in Figure 10 (a) and (b), the accuracy first increases and then drops as $\tau_{1}$ increases, which is maximized at $\tau_{1}=3$ on CiteULike and $\tau_{1}=4$ on Yelp. $\tau_{1}$ controls the strength of penalties on the pairs severely suffering from the collapse issue (i.e., highly correlated), the samples that are not highly correlated tend to be ignored when $\tau_{1}$ is set too large, while a too small $\tau_{1}$ fails to optimize the highly correlated pairs in an effective way (DirectSpec+ degenerates to DirectSpec when $\tau_{1}\rightarrow 0$). We take the relations among unobserved pairs into consideration via a polynomial graph filter $\sum_{l=0}^{L}\mathbf{\hat{A}}^{l}$ as reducing $\tau_{0}<\tau_{1}$ (i.e., the pairs showing closer relations are optimized with a lower intensity$<\tau_{1}$). In Figure 10 (c), DirectSpec+ achieves the best performance at $L=3$, incorporating neighborhood farther than $L>3$ results in degradation as they are too far away on the graph. In Figure 10 (d), the accuracy increases and then drops as reducing $\tau$, the maximum is reached at $\tau=3.0$ on Gowalla, demonstrating it effectiveness. We also observe that the improvement from adjusting $\tau_{0}$ is less significant on the other two datasets, which might be related to the data density, as users/items are less hierarchically distributed on denser data. Therefore, we can simple fix: $\tau_{0}=\tau_{1}$ for such datasets.
5.3.5. What Factors Affect DirectSpec’s Performance
As analyzed previously, DirectSpec’s performance is primarily influenced by the extent of collapse and the choice of encoders (i.e., the algorithm implemented on). The variation in improvement caused by different degrees of collapse is unavoidable as our goal is to eliminate the collapse issue. We can choose the encoder according to the data density: GNN-based methods perform better on sparse data while MF-based methods excel in dense data.
6. Related Work
6.1. Collaborative Filtering
@@
Collaborative Filtering (CF), a fundamental task for recommender systems, makes predictions based on user’s past behaviours. Early CF methods exploit users that share similar interests or items that tend to be interacted by similar users to infer user preference (Sarwar et al., 2001). Nowadays, model-based methods are prevalent in CF and recommender systems. The core idea is to map the high dimensional sparse recommendation data to a low-dimensional space, where users and items are characterized as learnable vectors and are optimized based on observed interactions (Koren et al., 2009). With the development of computer hardware and communication network, more information can be collected to analyze user behaviour with advanced algorithms. Besides user-item interactions, side information such as social relations (Jiang et al., 2012), reviews (Bao et al., 2014), geological information (Lian et al., 2014; Peng et al., 2018), and so on can help the algorithms better understand user taste and preference. On another line, early simple CF methods (Koren et al., 2009; Ning and Karypis, 2011) are replaced by more powerful algorithms such as neural network (Sedhain et al., 2015; He et al., 2017), attention mechanism (Chen et al., 2017; Kang and McAuley, 2018), transformer (Sun et al., 2019), etc. Recently, graph neural network (GNN) has also shown tremendous success in recommender systems (Wang et al., 2019; Peng et al., 2022b, 2024). By representing interactions as a bipartite graph, the core idea of GNN is to exploit higher-order neighbor connections to facilitate user/item representations (Kipf and Welling, 2017). However, it has been reported that multi-layer GNNs suffer from an over-smoothing issue that causes user/item representations to be indistinguishable (Peng et al., 2022a; Li et al., 2018). We showed that an embedding collapse issue resembling the over-smoothing exists on recommendation algorithms in general, and our proposed approach has also been shown effective tackling over-smoothing.
6.2. Collapse in Representation Learning
@@
Representation learning has attracted tremendous attention in various research fields (Mikolov et al., 2013; Hjelm et al., 2019), the goal of which is to represent data in an effective and efficient way to make it easier to extract useful information when building predictors. However, it has also been shown that the model tends to collapse where all inputs are mapped to the same constant vector, when only optimizing the model based on the positive pairs (Chen and He, 2021). This issue can be well alleviated by self-supervised learning and contrastive learning by exploiting negative samples in an effective way (Hjelm et al., 2019; Chen et al., 2020a). Due to the heavy computation cost, some research effort has been made to simplify the algorithms without explicitly sampling negative data (Chen and He, 2021; Grill et al., 2020; Zbontar et al., 2021; Zhang et al., 2023). Unfortunately, subsequent works show that a dimensional collapse cannot be ignored where the embedding vectors end up spanning a lower dimensional subspace of the whole embedding space (Hua et al., 2021; Jing et al., 2022). The collapse issue in representation learning shares similarities to some issues such as over-smoothing in GNN (Li et al., 2018), which inspires some researchers to tackle issues in GNN (Guo et al., 2023).
[…]
7. Conclusion and Limitation
@@
In this work, we showed that existing CF methods mostly suffer from an embedding collapse issue. Particularly, alignment of positive pairs is equivalent to a low pass filter, causing the representations of users and items to collapse to a constant vector. Exploiting negative pairs (i.e., negative sampling) can decelerate collapse by acting as a high pass filter to balance the embedding spectrum, while the effectiveness is only limited to certain loss functions which still results in an incomplete collapse, where embeddings are distributed along certain dimensions instead of making full use of all dimensions. To tackle this issue, we proposed a DirectSpec which directly optimizes the spectrum distribution to ensure that all dimensions can contribute to the user/item embeddings as equally as possible. We further pointed out that the key of high-quality representation lies in trade-off between alignment and spectrum balancing, and proposed an enhanced algorithm DirectSpec+, which can balance the spectrum more adaptively and individually. Moreover, we established the connection with SSL-based methods without explicit negative sampling such as Barlow-Twin, LogDet, and spectral contrastive loss by showing that they can be considered as special cases of DirectSpec. Finally, extensive experiments on three publicly available datasets demonstrated the effectiveness and efficiency of our proposed methods. Given the generality of our DirectSpec learning paradigm, we believe that our proposed method can effectively address collapse in representation learning more broadly.
We acknowledge that our proposed methods are still far from perfect, despite the efficiency and effectiveness. Firstly, balancing the embedding spectrum requires tuning hyperparameters such as $\alpha$ and $\tau$, reducing the utility. Furthermore, the effectiveness of some designs (e.g., $\tau_{0}$) is not significant and varies on datasets. We aim to address above limitations in the future work.
Acknowledgement
This paper is based on results obtained from the project, “Research and Development Project of the Enhanced Infrastructures for Post-5G Information and Communication Systems” (JPNP20017), commissioned by the New Energy and Industrial Technology Development Organization (NEDO).